OpenStudy (anonymous):

Graph the function. y = 1/5(3)^x

OpenStudy (anonymous):

answer choices

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@goformit100

OpenStudy (goformit100):

Sorry Sir I am weak at graphical approach of mathematics

OpenStudy (goformit100):

@sourwing

OpenStudy (anonymous):

@Destinymasha

OpenStudy (destinymasha):

Yea, I'm not too good with these either, sorry >.<

OpenStudy (anonymous):

uhhh.... graphing calculator is your best friend :D

OpenStudy (anonymous):

I don't have one

OpenStudy (sweetburger):

if u have an iphone or android download an emulator

OpenStudy (kainui):

I can help! Let's reason through each graph as to why it's either right or wrong.

OpenStudy (sweetburger):

anyways to solve this you can start by picking values for x and finding the corresponding y values

OpenStudy (kainui):

Often times we just need to know certain kinds of trends, not the exact graph at all. For exponential graphs, it's usually best to check where it is at x=0. Another good thing to notice is is it exponential growth or decay?

OpenStudy (kainui):

I think you're wrong @sweetburger As x increases, y increases.

OpenStudy (tkhunny):

It's strictly positive. Discard #2 and #4 y-intercept is 1/5. Discard #1 Not much left. \(\dfrac{1}{5}3^{2} = \dfrac{9}{5}\) \(\dfrac{1}{5}3^{3} = \dfrac{27}{5}\)

OpenStudy (kainui):

I'm pretty sure it's supposed to be: \[y=\frac{1}{5}3^x\]

OpenStudy (sweetburger):

well I read that wrong

OpenStudy (sweetburger):

ill see myself out

OpenStudy (anonymous):

yes you are correct @Kainui sorry about the confusion

OpenStudy (kainui):

@tkhunny is correct, but doesn't explain much. I'll just leave everyone else to figure this out because it's too crowded in here.

OpenStudy (anonymous):

so it is C??

OpenStudy (tkhunny):

There should not have been any confusion. The strictest application of Order of Operations conventions results in a proper understanding of the expression. 1/5(3)^x = (1/5)(3)^x = (1/5)(3^x) = \(\dfrac{1}{5}3^{x}\) There really is not much to explain. Read my post a few more times. That's ALL there is to it.

OpenStudy (kainui):

@tkhunny You're wrong. How did you determine the graph was strictly positive? To the uninitiated this isn't something people are born knowing. Most people have trouble understanding what negative or zero exponents mean. @wesniki23 Yes, the answer is C. But make sure you think about it and understand why because otherwise you'll have problems in the future. It's better to try to understand it now while we can still help you out. =)

OpenStudy (tkhunny):

\(1/5 > 0\) -- Can we assume people know that? \(3^{x}\) gets smaller and smaller with increasing negative values for x. It never turns negative. This only grows as x increases in the positive direction. There is some learning required for exponents. By the time you are graphing exponential functions, you should have it down pretty well. If you don't, I can usually blame a teacher or a curriculum.